if 2a+b+2c=0, then the family of lines ax+by+c=0 is concurrent at??

Grade:

1 Answers

Vikas TU
10 years ago

From 2a+b+2c=0

c = -(2a + b)/2 ..........................................(1)

Put (1) in ax+by+c=0

=> ax+by - a - b/2 = 0

=> a(x-1) + b(y - 1/2) = 0

{ comparing coefficients to zero we get,

a(x-1) = 0  or x = 1

and

b(y - 1/2) = 0 or y = 1/2

 

Hence family of lines ax+by+c=0 is concurrent at (1,1/2)

 

Plz Approve!

Plz Approve!

Plz Approve!